# Which Viewpoint Shows it Best? Language for Weakly Supervising View Selection in Multi-view Videos

### v2

@@ Abstract
Given a multi-view video, which viewpoint is most informative for a human observer? Existing methods rely on heuristics or expensive “best-view" supervision to answer this question, limiting their applicability. We propose a weakly supervised approach that leverages language accompanying an instructional multi-view video as a means to recover its most informative viewpoint(s). Our key hypothesis is that the more accurately an individual view can predict a view-agnostic text summary, the more informative it is. To put this into action, we propose a framework that uses the relative accuracy of view-dependent caption predictions as a proxy for best view pseudo-labels. Then, those pseudo-labels are used to train a view selector, together with an auxiliary camera pose predictor that enhances view-sensitivity. During inference, our model takes as input only a multi-view video—no language or camera poses—and returns the best viewpoint to watch at each timestep. On two challenging datasets comprised of diverse multi-camera setups and how-to activities, our model consistently outperforms state-of-the-art baselines, both with quantitative metrics and human evaluation. Project: http://vision.cs.utexas.edu/projects/which-view-shows-it-best.
@@ 1 Introduction
Figure 1: Main idea: given multi-view instructional videos, we aim to learn a view selection model that can identify the best view for seeing how to perform the activity shown in the videos, in the absence of best view labels. To achieve this, we compare each estimated view-dependent caption to the view-agnostic ground-truth video narration of the human activity, and use their respective accuracies as a proxy for view quality. These quality scores then serve as pseudo-labels for learning to select the most informative view. In this example, the 1st view most clearly shows all entities involved in the activity—the wheel and the person’s hands, and how they interact—and hence, produces a caption that best matches the ground-truth, making it a positive pseudo-label for view selection.
@@
To validate our hypothesis, we design a novel framework composed of two key elements: a best view pseudo-labeler, and a best view selector. The pseudo-labeler automatically generates best view pseudo-labels for a multi-view video during training, by using off-the-shelf video captioners [121, 64] to score and rank views on the basis of how well the predicted narration from a view matches the view-agnostic ground-truth narration. The selector takes a multi-view video as input, and predicts the best view labels. During training, the selector also solves an auxiliary task of predicting the relative camera pose between different views, to increase its view-sensitivity and improve its selection accuracy. At inference, our model requires as input only a multi-view video, but no language or camera poses.
We evaluate our method using two challenging multi-view instructional video datasets encompassing diverse activity scenarios and multi-camera setups, Ego-Exo4D [37] and LEMMA [56]. On both, our model outperforms multiple baselines and state-of-the-art methods for view selection on several automatic and human evaluation metrics. More broadly, our work offers a novel way for language to elicit the information content of video.
@@ 2 Related work
Automatic cinematography involves automatically choosing the best camera angles, positions, and zoom levels for human consumption of a video scene. Existing work explores camera control in virtual environments [30, 44, 45, 75], or addresses a narrow domain like lecture videos [31, 120, 100, 33, 46], social settings with multiple egocentric cameras [4], or panoramic 360 input videos [98, 97, 62, 18, 111]. Prior learning-based methods require manual labels to guide selection [98, 97, 62, 48, 13]. In contrast, we show how to train a view selector without manual labels, by exploiting the accuracy of predicted narrations from different views as a proxy for view quality.
@@ 3 Approach
Figure 2: (a) Our model uses language guidance to train a view-selector for multi-view instructional videos, such that the chosen views help best understand the shown activity. To do so, we first generate best view pseudo-labels during training by leveraging clip narrations, where each narration is a view-agnostic and detailed description of the activity. Specifically, given a training clip, we use off-the-shelf video captioners to predict a caption per view, score the views by comparing their captions to the ground-truth narration, and finally rank the views to generate a best view pseudo-label for the clip. Given the multi-view clip, our view classifier (bottom-left) encodes it with a visual encoder, and predicts a pseudo-label estimate. We also solve an auxiliary task of relative camera pose prediction (bottom-right) that increases the view sensitivity of the classifier. (b) Examples of predicted narrations, and the ranks and scores of the views per our pseudo-labeler, shown alongside ground-truth view-agnostic narrations. Note that at inference time, there is no ground truth narration, just the video input.
@@ 3.1 View selection task
Rather than assume any best-view labels on the training clips $\mathcal{V}_{m}$, we turn instead to narrations—human-provided descriptions of the activity in the video. Narrations are common in instructional video [123, 73, 36, 37, 56, 21], and for recent multi-view datasets [37, 56] they are specifically gathered in a view-agnostic manner: human annotators watch a collage of all views of a clip and write down a holistic description of the activity. See Fig. 2 (center).
@@ 3.3 Best view selector $S$
We use the pseudo-label set $\mathcal{B}$ from Sec. 3.2 to train our best view selector $S$, which must reason across all views to identify the best one. Simply using a view classifier as the selector is not enough in our setting, as our captioning-based labels can sometimes be insufficiently discriminative, i.e. , multiple views might be pseudo-labeled as the best view due to the very nature of the off-the-shelf video captioners’ training. They were originally designed to learn features that are predictive of the narration regardless of the view, and consequently, can end up collapsing all views into similar representations that are unable to capture the important nuances between them.
@@
To perform this classification for a viewpoint pair $(i,j)$, our pose predictor $P$ uses the fine-grained visual features $f_{i}$ and $f_{j}$ produced by our view classifier for viewpoints $i$ and $j$, and embeds them into more abstract representations $h^{P}_{i}$ and $h^{P}_{j}$ by using a feature projector $H^{P}$, such that $[h^{P}_{i},h^{P}_{j}]=[H^{P}(f_{i}),H^{P}(f_{j})]$. Whereas the fine-grained features $f$ help learn patch-level correspondences [8, 85], which are crucial for accurate pose prediction, the features $h$ act as a bridge between these detailed representations and the higher-level measure of relative pose. Finally, we concatenate [86] $h^{P}_{i}$ and $h^{P}_{j}$, and pass them to a pose classification head $C^{P}$, which computes their inter-feature correlation [10] and outputs a relative pose estimate $P_{i,j}$. Formally, $P_{i,j}=C^{P}([h^{P}_{i},h^{P}_{j}])$. The estimates are used in an auxiliary loss defined below. See Fig. 2(a) bottom-right.
@@ 4.1 Experimental setup
We evaluate our model on two multi-view instructional video datasets: Ego-Exo4D [37] and LEMMA [56] [^3] Note that other instructional video datasets like HowTo100M [73], CrossTask [124], and COIN [102] are single-view and thus not suitable.. Ego-Exo4D contains both physical (e.g. basketball, dancing) and procedural (e.g. cooking, bike repair) activities with each video containing 5 time-synced views—one is egocentric (ego) and recorded by the participant with a headworn camera, and the remaining views are exocentric (exo) and recorded with stationary cameras kept around the scene. LEMMA has videos of people performing household activities (e.g. making juice, watering plant), where each video has the participant’s ego view and an exo view from a fixed camera placed facing the activity. Both have narrations: the Ego-Exo4D annotators provide temporally dense written descriptions of the participants’ actions and relevant objects involved in the activity, while the LEMMA annotators describe action verbs and objects being interacted with using a pre-defined vocabulary. These two datasets let us evaluate different exo view setups (multiple in EgoExo-4D vs. single in LEMMA) and diverse activity scenarios (physical and procedural in Ego-Exo4D vs. household in LEMMA). Ego-Exo4D and LEMMA provide a total of 86 and 20 hours of video data, resulting in 648,665 and 63,538 clip-narration pairs, respectively. See Supp. for details.
@@
Through human evaluation, we assess two important model aspects: 1) pseudo-label quality, and 2) view selection performance. We conduct both assessments through pairwise comparison of views [20, 2, 34]. Given a view pair, the human judge can select either view or both, depending on if they prefer one over the other (i.e., a win for the preferred view, and a loss for the other), or find them equally informative (tie). Critically, we do not show the evaluators the ground-truth narrations. In other words, our study directly evaluates the human-preferred viewpoints—independent of narrations—and hence is unbiased by the fact our model leverages language during training. We obtain human judgments on 1) pseudo-label quality by pairing the best and the worst views per our pseudo-labeler (cf. Sec. 3.2), and 2) view prediction quality by pairing the views predicted by our and the best baseline’s predicted views. We decide the view order in each pair randomly. We do each study for both datasets with 10 participants and 70 randomly chosen test view pairs.
@@ 4.2 Results
Our model significantly outperforms all baselines across metrics, on both datasets. It shows our idea of pseudo-labeling by leveraging the view-dependent quality of predicted narrations is effective in practice. Furthermore, our model’s improvement over the heuristics illustrates that our language-guided training facilitates complex but essential reasoning about the interplay between human actors and interacting objects, more than what is possible with hand and object, or body pose detectors. Finally, our superior performance on both datasets underlines the efficacy of our design and its ability to generalize to different activity types—both physical and non-physical in Ego-Exo4D vs. household in LEMMA— and camera setups—single vs. multiple exo cameras in LEMMA and Ego-Exo4D, respectively. See Supp. for results on the single exo camera variant for Ego-Exo4D.
@@
View prediction
@@
Fig. 3 (left) shows our model’s success cases. Note how our model chooses views that clearly show all the essential elements of the actions, including the different entities involved and and their motion. E.g., in clip 2, our model chooses the ego view as it clearly shows the horizontal knife motion and vegetables on the cutting board, whereas in clip 3, it selects an exo view that best shows the joined hands and the sideways movement of the dancers. Thus, depending on the activity, our model adaptively chooses a view that accurately shows its important details. See Supp. video for more examples.
@@
See Supp. for additional ablations, and analyses of the view dependence of our visual features, our pseudo-labeler, and the impact of the rank of our selector’s sampled view on view selection performance.
@@ 5 Conclusion
We tackle view selection in multi-view instructional videos in the absence of best view labels. To that end, we design a novel framework composed of a best view pseudo-labeler that uses the view-dependent quality of estimates of video descriptions to automatically generate best view pseudo-labels, and a best view selector that given a video, produces a best view prediction. Our method significantly outperforms several state-of-the-art baselines on two challenging multi-view instructional video datasets. In future work, we will explore future best-view anticipation for improving the energy efficiency of multi-view instructional video capture setups.

### v4

# Which Viewpoint Shows it Best? Language for Weakly Supervising View Selection in Multi-view Instructional Videos

@@ Abstract
Given a multi-view video, which viewpoint is most informative for a human observer? Existing methods rely on heuristics or expensive “best-view" supervision to answer this question, limiting their applicability. We propose a weakly supervised approach that leverages language accompanying an instructional multi-view video as a means to recover its most informative viewpoint(s). Our key hypothesis is that the more accurately an individual view can predict a view-agnostic text summary, the more informative it is. To put this into action, we propose LangView, a framework that uses the relative accuracy of view-dependent caption predictions as a proxy for best view pseudo-labels. Then, those pseudo-labels are used to train a view selector, together with an auxiliary camera pose predictor that enhances view-sensitivity. During inference, our model takes as input only a multi-view video—no language or camera poses—and returns the best viewpoint to watch at each timestep. On two challenging datasets comprised of diverse multi-camera setups and how-to activities, our model consistently outperforms state-of-the-art baselines, both with quantitative metrics and human evaluation. Project: https://vision.cs.utexas.edu/projects/which-view-shows-it-best.
@@ 1 Introduction
Figure 1: LangView idea: given multi-view instructional videos, we aim to learn a view selection model that can identify the best view for seeing how to perform the activity shown in the videos, in the absence of best view labels. To achieve this, we compare each estimated view-dependent caption to the view-agnostic ground-truth video narration of the human activity, and use their respective accuracies as a proxy for view quality. These quality scores then serve as pseudo-labels for learning to select the most informative view. In this example, the 1st view most clearly shows all entities involved in the activity—the wheel and the person’s hands, and how they interact—and hence, produces a caption that best matches the ground-truth, making it a positive pseudo-label for view selection.
@@
To validate our hypothesis, we design a novel framework called LangView, which is composed of two key elements: a best view pseudo-labeler, and a best view selector. The pseudo-labeler automatically generates best view pseudo-labels for a multi-view video during training, by using off-the-shelf video captioners [122, 64] to score and rank views on the basis of how well the predicted narration from a view matches the view-agnostic ground-truth narration. The selector takes a multi-view video as input, and predicts the best view labels. During training, the selector also solves an auxiliary task of predicting the relative camera pose between different views, to increase its view-sensitivity and improve its selection accuracy. At inference, our model requires as input only a multi-view video, but no language or camera poses.
We evaluate LangView using two challenging multi-view instructional video datasets encompassing diverse activity scenarios and multi-camera setups, Ego-Exo4D [37] and LEMMA [56]. On both, our model outperforms multiple baselines and state-of-the-art methods for view selection on several automatic and human evaluation metrics. More broadly, our work offers a novel way for language to elicit the information content of video.
@@ 2 Related work
Automatic cinematography involves automatically choosing the best camera angles, positions, and zoom levels for human consumption of a video scene. Existing work explores camera control in virtual environments [30, 44, 45, 76], or addresses a narrow domain like lecture videos [31, 121, 101, 33, 46], social settings with multiple egocentric cameras [4], or panoramic 360 input videos [99, 98, 62, 18, 112]. Prior learning-based methods require manual labels to guide selection [99, 98, 62, 48, 13]. In contrast, we show how to train a view selector without manual labels, by exploiting the accuracy of predicted narrations from different views as a proxy for view quality. In continued complementary research, we explore how to leverage single-view instructional videos during training [72]
@@ 3 Approach
Figure 2: (a) Our model uses language guidance to train a view-selector for multi-view instructional videos, such that the chosen views help best understand the shown activity. To do so, we first generate best view pseudo-labels during training by leveraging clip narrations, where each narration is a view-agnostic and detailed description of the activity. Specifically, given a training clip, we use off-the-shelf video captioners to predict a caption per view, score the views by comparing their captions to the ground-truth narration, and finally rank the views to generate a best view pseudo-label for the clip. Given the multi-view clip, our view classifier (bottom-left) encodes it with a visual encoder, and predicts a pseudo-label estimate. We also solve an auxiliary task of relative camera pose prediction (bottom-right) that increases the view sensitivity of the classifier. (b) Examples of predicted narrations, and the ranks and scores of the views per our pseudo-labeler, shown alongside ground-truth view-agnostic narrations. “C" refers to the person who is performing the activity. Note that at inference time, there is no ground truth narration, just the video input.
@@ 3.1 View selection task
Rather than assume any best-view labels on the training clips $\mathcal{V}_{m}$, we turn instead to narrations—human-provided descriptions of the activity in the video. Narrations are common in instructional videos [124, 74, 36, 37, 56, 21], and for recent multi-view datasets [37, 56] they are specifically gathered in a view-agnostic manner: human annotators watch a collage of all views of a clip and write down a holistic description of the person’s actions and the objects involved. See Fig. 2 (center). While the narrations themselves are a form of annotation, they are more widely available, versatile, and scalable than specialized best-view labels, and hence provide a compelling source of weak supervision.
@@ 3.3 Best view selector $S$
We use the pseudo-label set $\mathcal{B}$ from Sec. 3.2 to train our best view selector $S$, which must reason across all views to identify the best one  [^3] Note that it is not possible to simply apply the pseudo-labeler at inference time, since it requires access to ground-truth narrations.. Simply using a view classifier as the selector is not enough in our setting, as our captioning-based labels can sometimes be insufficiently discriminative, i.e. , multiple views might be pseudo-labeled as the best view due to the very nature of the off-the-shelf video captioners’ training. They were originally designed to learn features that are predictive of the narration regardless of the view, and consequently, can end up collapsing all views into similar representations that are unable to capture the important nuances between them.
@@
To perform this classification for a viewpoint pair $(i,j)$, our pose predictor $P$ uses the fine-grained visual features $f_{i}$ and $f_{j}$ produced by our view classifier for viewpoints $i$ and $j$, and embeds them into more abstract representations $h^{P}_{i}$ and $h^{P}_{j}$ by using a feature projector $H^{P}$, such that $[h^{P}_{i},h^{P}_{j}]=[H^{P}(f_{i}),H^{P}(f_{j})]$. Whereas the fine-grained features $f$ help learn patch-level correspondences [8, 86], which are crucial for accurate pose prediction, the features $h$ act as a bridge between these detailed representations and the higher-level measure of relative pose. Finally, similar to VLocNet++ [87], we concatenate features $h^{P}_{i}$ and $h^{P}_{j}$, and pass them to a pose classification head $C^{P}$, which computes their inter-feature correlation [10] and outputs a relative pose estimate $P_{i,j}$. Formally, $P_{i,j}=C^{P}([h^{P}_{i},h^{P}_{j}])$. The estimates are used in an auxiliary loss defined below. See Fig. 2(a) bottom-right.
@@ 4.1 Experimental setup
We evaluate our model on two multi-view instructional video datasets: Ego-Exo4D [37] and LEMMA [56] [^4] Note that other instructional video datasets like HowTo100M [74], CrossTask [125], and COIN [103] are single-view and thus not suitable.. Ego-Exo4D contains both physical (e.g. basketball, dancing) and procedural (e.g. cooking, bike repair) activities with each video containing 5 time-synced views—one is egocentric (ego) and recorded by the participant with a headworn camera, and the remaining views are exocentric (exo) and recorded with stationary cameras kept around the scene. LEMMA has videos of people performing household activities (e.g. making juice, watering plant), where each video has the participant’s ego view and an exo view from a fixed camera placed facing the activity. Both have narrations: the Ego-Exo4D annotators provide temporally dense written descriptions of the participants’ actions and relevant objects involved in the activity, while the LEMMA annotators describe action verbs and objects being interacted with using a pre-defined vocabulary. These two datasets let us evaluate different exo view setups (multiple in EgoExo-4D vs. single in LEMMA) and diverse activity scenarios (physical and procedural in Ego-Exo4D vs. household in LEMMA). Ego-Exo4D and LEMMA provide a total of 86 and 20 hours of video data, resulting in 648,665 and 63,538 clip-narration pairs, respectively. See Supp. for details, including our clip segmentation strategy.
@@
Through human evaluation, we assess two important model aspects: 1) pseudo-label quality, and 2) view selection performance. We conduct both assessments through pairwise comparison of views, which reduces the cognitive load on human judges and increases their reliability [20, 2, 34]. Given a view pair, the human judge can select either view or both, depending on if they prefer one over the other specifically for the purpose of activity understanding (i.e., a win for the preferred view, and a loss for the other), or find them equally informative (tie). Critically, we do not show the evaluators the ground-truth narrations. In other words, our study directly evaluates the human-preferred viewpoints—independent of narrations—and hence is unbiased by the fact our model leverages language during training. We obtain human judgments on 1) pseudo-label quality by pairing the best and the worst views per our pseudo-labeler (cf. Sec. 3.2), and 2) view prediction quality by pairing the views predicted by our and the best baseline’s predicted views. We decide the view order in each pair randomly. We do each study for both datasets with 10 participants and 70 randomly chosen test view pairs. Our inter-evaluator agreement rate is 78.5%.
@@ 4.2 Results
Our model significantly outperforms all baselines across metrics, on both datasets. It shows our idea of pseudo-labeling by leveraging the view-dependent quality of predicted narrations is effective in practice. Furthermore, our model’s improvement over the heuristics illustrates that our language-guided training facilitates complex but essential reasoning about the interplay between human actors and interacting objects, more than what is possible with hand and object, or body pose detectors. Finally, our superior performance on both datasets underlines the efficacy of our design and its ability to generalize to different activity types—both physical and non-physical in Ego-Exo4D vs. household in LEMMA— and camera setups—single vs. multiple exo cameras in LEMMA and Ego-Exo4D, respectively. See Supp. for results on the single exo camera and 3-fold evaluation with Ego-Exo4D.
@@
View prediction: ours vs best baselines
@@
Fig. 3 (left) shows some success cases. Note how our model chooses views that clearly show all the essential elements of the actions, including the different entities involved and and their motion. E.g., in clip 2, our model chooses the ego view as it clearly shows the horizontal knife motion and vegetables on the cutting board, whereas in clip 3, it selects an exo view that best shows the joined hands and the sideways movement of the dancers. Thus, depending on the activity, our model adaptively chooses a view that accurately shows its important details. See Supp. video for more examples.
@@
See Supp. for additional ablations, and analyses of the view dependence of our visual features, our pseudo-labeler, and the impact of the rank of our selector’s sampled view on view selection performance, and our model’s attention maps.
@@ 5 Conclusion
We tackle view selection in multi-view instructional videos in the absence of best view labels. To that end, we design a novel framework composed of a best view pseudo-labeler that uses the view-dependent quality of estimates of video descriptions to automatically generate best view pseudo-labels, and a best view selector that given a video, produces a best view prediction. Our method significantly outperforms several state-of-the-art baselines on two challenging multi-view instructional video datasets. In future work, we will explore future best-view anticipation for improving the energy efficiency of multi-view instructional video capture setups. Acknowledgements: UT Austin is supported in part by the IFML NSF AI Institute. KG is paid as a research scientist by Meta, and SM was a visiting researcher at the same when this work was done.